7x + 9y = -27
-2x - 7y = 21
Solve using either elim. or sub.

Answer by mananth(16949)

7x + 9y = -27
-2x - 7y = 21

2(7x+9y)+7(-2x-7y)= 2*-27 + 7*21
14x+18y-14x-49y=-54+147
-31y=93
y=-3
-2x-7*(-3)=21
x=0